A body of mass 40kg is given an acceleration of 10 meter per seconds on a horizontal ground for which m=0.5. Calculate the force required to accelerate to the body.( take g=10 meter per seconds )
1
Expert's answer
2021-08-16T08:37:33-0400

According to Newton's second law:


"ma=F-f,\\
ma=F-\mu mg,\\\space\\
F=ma+\mu mg=m(a+\mu g)=596\text N."

Actually, the force of friction is


"f=\mu N,"

where N - normal force. But since the body is on a horizontal surface, we can state that


"N=mg."